# COCO: Cognitive Operating System with Continuous Oversight for Multi-Agent Workflow Reliability

### v1

@@ Abstract
Large-scale multi-agent workflows exhibit inherent vulnerability to error propagation and quality degradation, where downstream agents compound upstream failures without corrective mechanisms. We introduce COCO (Cognitive Operating System with Continuous Oversight), a theoretically-grounded framework that implements asynchronous self-monitoring and adaptive error correction in multi-agent driven systems. COCO addresses the fundamental trade-off between quality assurance and computational efficiency through a novel decoupled architecture that separates error detection from the critical execution path, achieving $O(1)$ monitoring overhead relative to workflow complexity. COCO employs three key algorithmic innovations to address systematic and stochastic errors: (1) Contextual Rollback Mechanism - a stateful restart protocol that preserves execution history and error diagnostics, enabling informed re-computation rather than naive retry; (2) Bidirectional Reflection Protocol - a mutual validation system between monitoring and execution modules that prevents oscillatory behavior and ensures convergence; (3) Heterogeneous Cross-Validation - leveraging model diversity to detect systematic biases and hallucinations through ensemble disagreement metrics. Extensive experiments on benchmark multi-agent tasks demonstrate 6.5% average performance improvement, establishing new state-of-the-art for autonomous workflow reliability.
Introduction
The paradigm of compositional intelligence through multi-agent systems has emerged as a cornerstone of modern artificial intelligence(Wu 2023), enabling unprecedented scalability and modularity in complex reasoning tasks(Qian 2025). These architectures decompose intricate problems into specialized computational modules that collaborate through sequential and parallel information flows, achieving remarkable performance across domains ranging from scientific discovery to autonomous decision-making(Li 2025). However, this compositional approach introduces a fundamental trade-off: while modularity enhances interpretability and scalability, the inherent interdependencies between agents create pathways for systematic error amplification that can catastrophically degrade system performance(Cemri 2025).
@@ Introduction
The Error Propagation Problem. Consider a multi-agent workflow $\mathcal{W}=\{A_{1},A_{2},\ldots,A_{n}\}$ with where each agent $A_{i}$ processes intermediate representations $z_{i-1}$ to produce outputs $z_{i}=A_{i}(z_{i-1};\theta_{i})$. The fundamental vulnerability arises from the compounding nature of approximation errors: even small perturbations $\epsilon_{i}$ in upstream outputs can lead to exponential degradation in final quality(Gan 2025; Peng 2025),formalized as $\mathbb{E}[\|z_{n}-z_{n}^{*}\|]\leq\prod_{i=1}^{n}(1+\epsilon_{i})$ where $z_{n}^{*}$ denotes the optimal output. This epistemic error cascading destabilizes long pipelines, and error grows exponentially with pipeline depth.(Caselli 2015)
Figure 1: When agents make mistakes, COCO steps in to help—by catching errors early and enabling self-repair.
Limitations of Existing Approaches. Existing fault-tolerance strategies in multi-agent systems struggle to meet low-latency response, self-repairability demands. Circuit-breaker patterns intercept pathological outputs during inference but lack task-level semantics, making intelligent re-planning or self-repair infeasible (Zou 2024). Gatekeeping mechanisms like LLMGuard (Goyal 2024) proactively block harmful or hallucinated content, yet rely on external control and offer no recovery path. Synchronous validation frameworks such as SelfCheck and Self-Refine improve reasoning accuracy but introduce extra latency, making them impractical for latency-sensitive applications (Manakul 2023; Madaan 2023).
COCO Framework. COCO (Cognitive Operating System with Continuous Oversight) is thus proposed to effectively resolve both systematic and stochastic errors in multi-agent workflows while maintaining minimal computational overhead. The core of COCO lies the Contextual Rollback Mechanism (CRM), which implements intelligent state recovery through rollback optimization that preserves execution history and error diagnostics, enabling informed re-computation rather than naive retry. And COCO employs Bidirectional Reflection Protocol (BRP) combining Reflection and Chain-of-Thought (CoT) approaches to establish consensus-driven error correction with formal convergence guarantees, preventing correction oscillations through symmetric evaluation between execution and monitoring modules. Finally, Heterogeneous Cross-Validation (HCV) leverages ensemble disagreement analysis across diverse models to detect systematic biases and hallucinations through information-theoretic measures, providing bias-resistant detection capabilities. Contributions of the proposed approach can be summarized as:
@@
  • The proposed approach redefines error handling in multi-agent systems through asynchronous decoupled monitoring. Unlike existing approaches that create bottlenecks through synchronous validation or lack intelligent recovery capabilities, COCO maintains comprehensive error detection with low computing cost overhead.
  • Our framework employs three key algorithmic innovations: (1) CRM: Preserves execution snapshots and augments prompts to guide the model away from previous failure paths during rollback.; (2)BRP: Sends error feedback to the execution node, which reflects on the issue and reconstructs its output based on the feedback and internal reasoning.; and (3) HCV: Leverages disagreement among models of different architectures to improve error detection and system robustness.
  • The proposed approach provides a complete platform with annotation-based integration that requires minimal changes to existing workflows, enabling seamless deployment across multi-agent systems through native time-travel mechanisms rather than custom state management infrastructure.
  • Experiments across four diverse benchmarks demonstrate that COCO bridges 53.9% of the performance gap between baseline Mimas-LLM driven systems and Titan-scale model, achieving 95.1% of state-of-the-art performance with 30× parameter reduction while delivering consistent 6.5% improvements over strong baselines.
Related Work
LLM-driven Agent. Large language models have progressed from single-turn generation to tool-augmented multi-step reasoning (Wei 2022; Yao 2023; Schick 2023). Those steps are now auto-orchestrated as graph workflows (Wu 2025), and further extended into multi-agent systems where heterogeneous models share subtasks (Sheng 2023; Chen 2024; Wang 2025; Islam 2024).
Collaborative Model Teams. In order to further enhance model capabilities and inference efficiency, a series of structured mechanisms based on collaborative modeling have recently emerged. Dynamic agent pools select the best model per token at decode time, yielding 5–10 point boosts in human evaluation and allowing graceful scaling as new experts are added (Chakraborty 2024). Fixed-role pipelines split retrieval, planning, generation, and debugging across different models, raising contest coding pass rates by roughly 30% under the same compute budget (Islam 2024). cross-team peer reviews in AgentVerse and AutoGen (Chen 2024; Sheng 2023) continually refine answers and enforce safety.
Error Snowballing. Latest empirical evidence shows that when a high-error-rate agent appears in a multi-agent or graph-structured workflow, the overall accuracy drops rapidly with interaction rounds; for certain linear structures the decline can exceed 30% (Huang 2025); The BEHAVIORCHAIN benchmark similarly observes that a first-step error is amplified layer by layer in multi-round cooperation (Zhou 2025). This “snowballing” phenomenon has been quantitatively characterized by an information-theoretic framework, proving that the first error weakens the global reasoning success rate exponentially (Gan 2025). In addition, UProp decomposes errors into intrinsic and extrinsic components from an uncertainty perspective, revealing their propagation rules in multi-step chains (Tan and Chen 2025), Counterfactual analysis further shows that precisely repairing only a few key nodes can halve the system failure rate  (Chen 2025). Together these researches indicate that, in collaborative networks, early errors are continuously amplified by upstream and downstream agents and must be curbed at the source.
Figure 2: COCO system architecture. The framework consists of three core components: (1) the Decorator Binding Layer attaches non-intrusive hooks to agent nodes for asynchronous monitoring; (2) the Monitoring Engine uses large language models to evaluate outputs and generate structured diagnostics; and (3) the Correction Engine performs context-aware rollback via CRM, enhanced with BRP and HCV to improve correction stability and robustness.
Self-Auditing Models. Self-monitoring further cuts hallucinations: SelfCheckGPT runs multi-sample consistency checks to flag internal contradictions and factual drift (Manakul 2023); Self-Refine iteratively loops generate–critique–rewrite, delivering an average 20% improvement across tasks (Madaan 2023); Chain-of-Verification lets the model draft its own Q&A pairs before revision, lowering open-domain QA hallucinations by 26% (Dhuliawala 2024).
Runtime Guardrails. To halt the snowballing of errors in workflow, three runtime guardrails for multi-agent systems are commonly employed: Crowdsourcing-style parallel validation with voting splits each sub-task into multiple candidate paths and lets verification agents vote, intercepting anomalies at the earliest stage (Grunde-McLaughlin 2024; Sun 2024). Gatekeepers such as LLMGuard employ multiple detectors to audit each input and output, blocking harmful or hallucinated content before it reaches the next node (Goyal 2024). Chaos engineering randomly drops, delays, or tampers with messages during runtime, stress-testing the system and validating rollback as well as redundancy strategies (Owotogbe 2025).
Methodology
Multi-agent AI systems face a fundamental reliability challenge: errors propagate cascadingly through interconnected workflows, where each agent’s mistakes amplify downstream failures. Traditional quality assurance approaches operate synchronously, creating performance bottlenecks that scale linearly with workflow complexity, while simple retry mechanisms lack the contextual awareness necessary for intelligent recovery. This creates an inherent tension between quality guarantees and computational efficiency—existing systems must choose between thorough validation and real-time performance.
COCO addresses these fundamental limitations through a novel approach that leverages native graph-based workflow capabilities for intelligent monitoring and automatic correction. Unlike conventional approaches that require custom state management systems, COCO directly utilizes existing checkpoint and time-travel functionality within workflow frameworks, achieving seamless integration with existing multi-agent systems. The framework implements asynchronous monitoring that operates in parallel with workflow execution, maintaining quality guarantees without sacrificing performance through O(1) monitoring overhead relative to workflow depth.
Problem Definition
@@
Context Loss During Recovery: Traditional retry mechanisms discard valuable execution context when errors occur, causing agents to repeat the same mistakes or introduce new errors during recovery attempts.
@@
Correction Oscillations: Naive error correction systems often generate conflicting fixes that prevent convergence, as competing correction mechanisms may introduce contradictory changes that fail to stabilize.
@@
Performance-Quality Tradeoff: Conventional validation approaches introduce significant computational overhead that scales linearly or worse with workflow complexity, forcing system designers to compromise between thorough validation and real-time performance.
Error Correction Components
@@
COCO leverages the complementary strengths of its three core mechanisms to address these challenges: Contextual Rollback Mechanism, Bidirectional Reflection Protocol and Heterogeneous Cross-Validation.
Contextual Rollback Mechanism
The Contextual Rollback Mechanism (CRM) addresses error repetition caused by context loss in traditional rollback strategies. Whereas conventional systems discard execution state on failure, CRM preserves structured snapshots for each workflow node—including agent state, prompt history, reasoning traces, diagnostics, and timestamps—enabling context-aware restarts. When abnormal output is detected, the monitor triggers a rollback where CRM: First extracts relevant context and error analysis from execution history. Then generates an augmented prompt incorporating error information. Next, introduces semantic perturbations to encourage the model to explore alternative reasoning paths. Finally, preserves critical context to prevent the system from repeating the same mistakes
By maintaining this contextual information, CRM reduces the risk of falling into locally convergent failure modes while providing agents with the information needed to correct their errors intelligently.
Bidirectional Reflection Protocol
The Bidirectional Reflection Protocol (BRP) implements a symmetric evaluation mechanism between execution and monitoring modules, facilitating stable error negotiation. BRP combines two complementary processes:
Reflection: When detecting potential flaws, the monitor provides structured feedback—including error type, diagnostic rationale, and contextual trace—to the execution node, highlighting specific errors and suggesting correction strategies.
Chain-of-Thought: The execution node conducts self-reflection, integrating the original input, reasoning history, and monitor critique to assess feedback validity, enabling Chain-of-Thought reasoning for improved outputs.
This bidirectional interaction constitutes a form of self-alignment, where the system autonomously evaluates and corrects its reasoning. Unlike unidirectional monitoring approaches where monitors simply override execution outputs, BRP establishes a cooperative negotiation that reduces correction oscillations and improves convergence stability.
Heterogeneous Cross-Validation
@@
Overall Architecture

### v2

@@ Abstract
[…]
1 Introduction
The transition toward compositional intelligence via multi-agent systems has become a defining characteristic of modern AI Wu (2023). By decomposing high-dimensional reasoning tasks into specialized, collaborating modules, these architectures achieve unprecedented scalability and adaptability Qian et al. (2025). However, this modularity introduces a critical structural vulnerability: the sequential interdependencies required for collaboration create pathways for error amplification. Consequently, minor upstream hallucinations can cascade through the workflow, catastrophically compromising global system integrity Cemri et al. (2025).
@@ 1 Introduction
The Error Propagation Problem. For a formalized multi-agent workflow $W=\{A_{1},A_{2},\ldots,A_{n}\}$, where each agent $A_{i}$ transforms an intermediate representation $z_{i-1}$ into an output $z_{i}=A_{i}(z_{i-1};\theta_{i})$. A critical vulnerability in this sequential dependency is the compounding nature of approximation errors. Small perturbations $\epsilon_{i}$ at early stages do not merely persist; they act as multiplicative factors in downstream deviations Gan et al. (2025); Peng et al. (2025). This phenomenon is bounded by $\mathbb{E}[\|z_{n}-z_{n}^{*}\|]\leq\prod_{i=1}^{n}(1+\epsilon_{i})$ where $z_{n}^{*}$ denotes the optimal output trajectory. We term this epistemic error cascading, a mechanism where stochastic noise accumulates exponentially with pipeline depth, inevitably destabilizing complex reasoning tasks  Caselli et al. (2015).
Figure 1: By intercepting intermediate execution errors through asynchronous monitoring, COCO initiates a self-repair loop that prevents the downstream propagation of inaccuracies.
Limitations of Existing Approaches. Current fault-tolerance strategies fail to reconcile the tension between low-latency inference and autonomous self-repair. One category of solutions, including circuit-breaker patterns Zou et al. (2024) and gatekeeping mechanisms like LLMGuard Goyal et al. (2024), excels at interception but suffers from semantic unawareness. These systems function as terminal blockers: they can halt pathological outputs but lack the contextual understanding required to initiate intelligent recovery. Conversely, synchronous validation frameworks such as SelfCheck and Self-Refine Manakul et al. (2023); Madaan et al. (2023) improve reasoning fidelity but introduce prohibitive latency overheads. By forcing sequential verification steps, they create execution bottlenecks that render them impractical for scalable, latency-sensitive workflows.
COCO Framework. To resolve the fundamental tension between workflow reliability and computational efficiency, we propose COCO (Cognitive Operating System with Continuous Oversight). COCO departs from synchronous validation paradigms by adopting a decoupled asynchronous architecture. This design effectively isolates the Monitoring Engine from the critical execution path, ensuring $O(1)$ overhead relative to workflow complexity.
To address the specific failure modes of error amplification, oscillation, and bias, COCO orchestrates three synergistic mechanisms: (1) The Contextual Rollback Mechanism (CRM), which replaces naive retries with stateful recovery. By preserving execution snapshots ($\Psi_{i}$) and injecting diagnostic patches ($\Delta_{i}$), CRM enables informed re-computation that halts epistemic error cascading. (2) The Bidirectional Reflection Protocol (BRP), which mitigates control-loop instability. By establishing a symmetric consensus protocol between execution and monitoring nodes, BRP prevents correction oscillations and ensures convergence. (3) Heterogeneous Cross-Validation (HCV), which targets systematic bias. By leveraging information-theoretic disagreement across diverse model architectures, HCV detects subtle hallucinations that escape homogenous self-checks.
The main contributions are summarized as follows:
@@
  • Decoupled Asynchronous Oversight: We redefine error handling in multi-agent systems via a non-intrusive architecture that isolates oversight from the execution path. Unlike synchronous validators that create linear latency bottlenecks, COCO maintains comprehensive error detection with minimal computational overhead.
  • Algorithmic Novelty: We introduce a tripartite error-correction stack comprising CRM for history-aware state restoration instead of naive retries, BRP for achieving stable actor-critic consensus, and HCV for bias-resistant detection through heterogeneous ensemble disagreement analysis.
  • Production-Ready Integration: COCO provides a complete annotation-based platform that seamlessly integrates with native checkpointing mechanisms, supported by an automated configuration engine that optimizes monitoring strategies via empirical sampling.
[…]
2 Related Work
From Isolated Generation to Multi-Agent Workflows. The trajectory of large language models (LLMs) has shifted from isolated generation to the paradigm of compositional intelligence. Early methodologies enhanced individual reasoning capabilities through Chain-of-Thought (CoT) and ReAct paradigms, enabling models to decompose complex queries into executable steps Wei et al. (2023); Yao et al. (2023). Recent frameworks have scaled this capability into Multi-Agent Systems (MAS), where heterogeneous models are auto-orchestrated into graph-based workflows Zhang et al. (2025). Architectures such as AutoGen Wu et al. (2024) and AgentVerse Chen et al. (2023) distribute subtasks across specialized roles, while structured mechanisms like dynamic agent pools Chakraborty et al. (2025) and fixed-role pipelines Islam et al. (2024) optimize inference efficiency. However, this transition from single-turn inference to sequential, interdependent collaboration inherently increases the depth of the execution graph. As formalized before, this expanded interaction surface creates the structural conditions necessary for epistemic error cascading.
Dynamics of Error Propagation. Recent empirical studies validate the ”snowballing” phenomenon in epistemic error cascading. When a high-error-rate agent appears in a graph-structured workflow, global accuracy degrades exponentially with interaction rounds. In linear structures, this decline can exceed 30% tse Huang et al. (2025). The BEHAVIORCHAIN benchmark similarly observes that initial approximation errors are amplified layer-by-layer during multi-round cooperation Li et al. (2025).
This phenomenon has been analyzed through information-theoretic frameworks Gan et al. (2025) and uncertainty decomposition (UProp) Duan et al. (2026), which reveal that early-stage hallucinations function as multiplicative noise. Crucially, counterfactual analysis suggests that precisely repairing key nodes can halve system failure rates Chen et al. (2025), providing the theoretical motivation for the Contextual Rollback Mechanism employed in the proposed framework.
Auditing and Guardrail Paradigms. Current approaches to mitigating these errors fall into two primary categories, each with distinct limitations regarding the latency-reliability trade-off:
1) Synchronous Self-Correction: Methods such as SelfCheckGPT Manakul et al. (2023) and Self-Refine Madaan et al. (2023) employ iterative generate–critique–rewrite loops to reduce hallucinations. While effective—delivering average improvements of 20%—these approaches are inherently synchronous, introducing linear latency penalties that render them impractical for real-time applications.
2) Runtime Guardrails: To ensure safety without the latency of self-correction, systems employ external validation. Strategies include crowdsourcing-style parallel voting Grunde-McLaughlin et al. (2025) and deterministic gatekeepers like LLMGuard Goyal et al. (2024). While these mechanisms effectively intercept pathological outputs, they function as terminal blockers with limited semantic understanding, often lacking the context required for the intelligent state recovery that the proposed framework provides.
Figure 2: Architectural overview of the COCO framework. The system is composed of three decoupled layers: (1) The Decorator Binding Layer facilitates non-intrusive, asynchronous oversight via agent-level hooks; (2) The Monitoring Engine leverages LLMs to synthesize structured diagnostic reports from execution outputs; and (3) The Correction Engine orchestrates state restoration, integrating CRM for contextual rollback with BRP and HCV to guarantee correction stability.
3 Methodology
Multi-agent workflows are inherently vulnerable to error cascading, yet current mitigation strategies face a critical dichotomy: synchronous validation imposes linear latency penalties ($O(n)$), while stateless retries lack the semantic context for recovery. COCO reconciles this tension by leveraging native graph-based checkpointing to implement asynchronous decoupled monitoring. Unlike invasive state-management solutions, COCO integrates seamlessly with existing frameworks, utilizing parallel oversight to enforce quality guarantees with constant $\mathcal{O}(1)$ overhead relative to workflow depth.
3.1 Problem Definition
@@
Context Loss During Recovery: Traditional retry mechanisms discard the execution state $S_{i}$ upon failure. This loss of context causes agents to repeat identical mistakes or introduce new errors during recovery attempts due to a lack of informed diagnostic history.
@@
Correction Oscillations: Naive error correction systems often generate conflicting fixes that prevent convergence. Without a symmetric evaluation protocol like BRP, competing correction mechanisms may introduce contradictory changes that fail to stabilize the reasoning path.
@@
Performance-Quality Tradeoff: Conventional validation approaches introduce significant computational overhead that scales linearly, $O(n)$, with workflow depth. This forces system designers to compromise between thorough validation and real-time performance.
3.2 Error Correction Components
@@
COCO leverages the complementary strengths of its three core mechanisms to address these challenges: Contextual Rollback Mechanism (CRM), Bidirectional Reflection Protocol (BRP), and Heterogeneous Cross-Validation (HCV).
3.2.1 Contextual Rollback Mechanism
To mitigate the stochastic degeneracy inherent in stateless retry heuristics, CRM implements a history-aware state recovery protocol. Unlike naive approaches that discard the execution context upon failure, CRM persists a structured checkpoint $\Psi_{i}=\langle s_{i},P_{i}\rangle$ for each agent $A_{i}$, comprising the pre-execution state $s_{i}$ and the original instruction set $P_{i}$.
Trigger and Formulation. Upon the detection of a high-confidence failure—formalized as the condition $\mathbb{I}(h_{i}=1)\land(c_{i}\geq\alpha)$—CRM interrupts the downstream propagation and reverts the workflow to state $s_{i}$. Crucially, it constructs a Correction Patch $\Delta_{i}$ to condition the re-generation:
$\Delta_{i}=\langle\delta_{i},\gamma_{i},z_{i}^{fail}\rangle$ (1)
where $\delta_{i}$ represents the diagnostic rationale, $\gamma_{i}$ denotes the corrective constraints, and $z_{i}^{fail}$ serves as a negative sample.
Informed Re-computation. The re-execution mechanism is modeled as a counterfactual generation process:
$z_{i}^{\prime}=A_{i}(s_{i},P_{i}^{;}\theta_{i})\quad\text{s.t.}\quad P_{i}^{=}\mathcal{F}{aug}(P_{i},\Delta_{i})$ (2)
Here, the augmentation function $\mathcal{F}{aug}$ integrates the patch $\Delta_{i}$ into the prompt manifold. By explicitly incorporating the failure mode $z_{i}^{fail}$ and the diagnostic $\delta_{i}$, CRM effectively modifies the optimization landscape, penalizing the previously taken path. This transforms the recovery process from a random resampling event into a guided local search, ensuring that $z_{i}^{\prime}$ diverges from locally convergent failure modes and effectively curbing epistemic error cascading.
3.2.2 Bidirectional Reflection Protocol
To address the limitations of unidirectional validation, BRP implements a Dual-Phase Rectification Topology. Unlike iterative feedback systems that introduce unbounded latency, BRP operates as a strictly bounded, single-turn consensus mechanism. This design guarantees system stability by enforcing a hard break in the feedback loop, relying on the depth of reasoning (Chain-of-Thought) rather than the frequency of iteration to ensure quality. The protocol executes in two deterministic phases:
Phase 1: Diagnostic Projection (Monitor Side). Upon detecting an anomaly, the monitoring engine $\mathcal{M}$ does not merely reject the output but projects a structured diagnostic vector $\mathbf{R}_{i}$ into the agent’s context window. Formally, given the initial faulty output $z_{i}^{fail}$ and agent state $s_{i}$, the monitor computes:
$\mathbf{R}_{i}=\mathcal{M}(z_{i}^{fail},\Psi_{i})=\langle h_{i},c_{i},\tau_{i},\delta_{i},\gamma_{i}\rangle$ (3)
where $\Psi_{i}$ denotes structured execution snapshots, $h_{i}$ indicates the presence of errors, $c_{i}$ means error type, $\tau_{i}$ identifies the error category, $\delta_{i}$ provides the diagnostic rationale (the ”Why”), and $\gamma_{i}$ offers directive constraints (the ”How”) to guide the correction.
Phase 2: Reflective Reconstruction (Agent Side). Instead of simply retrying, the execution agent $A_{i}$ enters a Reflective Chain-of-Thought (CoT) mode. It consumes $\mathbf{R}_{i}$ to strictly condition its re-generation process. The corrected output $z_{i}^{\prime}$ is generated as:
$z_{i}^{\prime}=A_{i}(s_{i},z_{i}^{fail},\mathbf{R}i;\theta{cot})$ (4)
In this phase, the agent explicitly articulates an internal reasoning trace that aligns its intent with the constraints provided in $\gamma_{i}$ before producing the final result.
Stability and Overhead Analysis. This rectification architecture guarantees both structural stability and computational efficiency by enforcing a strict Directed Acyclic Graph (DAG) topology in the correction logic. By structurally severing the feedback edge after the reconstruction phase, BRP inherently eliminates the possibility of control-loop oscillation while ensuring bounded overhead with a deterministic $O(1)$ inference cost, thereby preventing the latency spikes characteristic of indeterminate iterative refinement.
3.2.3 Heterogeneous Cross-Validation
@@
Algorithm 1 Asynchronous Oversight and Contextual Recovery via COCO
Input: Workflow $\mathcal{G}(\mathcal{V},\mathcal{E})$, Input query $x$, Configuration $\mathcal{D}$
Output: Final verified output $z_{n}$
1:Phase 1: Configuration Overlay Initialization
2:if $\mathcal{D}$ is NULL then
3:  $\mathcal{D}\leftarrow\text{SampleAndTest}(\mathcal{G},\text{tasks})$ {Identify high-risk nodes by testing}
4:end if
[…]
6:Phase 2: Execution & Asynchronous Monitoring
7:for each agent node $A_{i}\in\text{Path}(\mathcal{G})$ do
8:  $z_{i}\leftarrow A_{i}(s_{i},P_{i};\theta_{exe})$ {Execute original agent task}
9:  if $A_{i}\in\tilde{\mathcal{V}}$ then
10:   Async Trigger:{Spawn parallel monitor thread}
11:   $\mathbf{R}_{i}=\langle h_{i},c_{i},\tau_{i},\delta_{i},\gamma_{i}\rangle\leftarrow\mathcal{M}(z_{i},\Psi_{i})$ {Generate diagnostics}
12:   if $h_{i}$ is True AND $c_{i}>\alpha$ then
13:    Phase 3: Context-Aware Recovery (CRM + BRP)
14:    $\Delta_{i}\leftarrow\text{GenPatch}(\Psi_{i},\mathbf{R}_{i})$ {Patch Generation via context and diagnostics}
15:    $P_{i}^{*}\leftarrow\mathcal{F}_{aug}(P_{i},\Delta_{i})$ {Prompt augmentation via the patch}
16:    $z_{i}^{\prime}\leftarrow A_{i}(s_{i},P_{i}^{*})$ {Reflective correction via BRP and CoT}
17:    Update $z_{i}\leftarrow z_{i}^{\prime}$ and re-route downstream execution
18:   end if
19:  end if
20:end for
21:return $z_{n}$
3.3 Overall Architecture